# Presenting a STEM Ways of Thinking Framework for Engineering Design-based Physics Problems

### v3

@@ 3.7 Implications of our Framework
In spite of the aforementioned limitations, our WoT4EDP framework has significant implications for teaching and learning in STEM education, as well as educational research. By explicitly identifying and characterizing different forms of thinking—design, science, mathematical, metacognitive, and computational—this framework provides educators with a structured yet flexible approach to supporting and analyzing student thinking during engineering design-based tasks. For instance, instructors can use the framework to design activities that intentionally foster these thinking skills or assess student work with a focus on how well students integrate design and science thinking. The framework also offers guidance (see ‘Enhancement approaches’ in Table 2) for scaffolding learning [78, 18], enabling educators to better identify gaps in student understanding and adapt instruction accordingly.
@@ 4 WoT Frameworks - A Comparative Study
FAWTEER emphasizes four specific ways of thinking: Futures, Values, Systems, and Strategic thinking. It conceptualizes Ways of Thinking as a framework that can be practically applied to drive innovation and enhance researchers’ capacity to effect systemic change. Notably, Dalal et al. acknowledge that other ways of thinking, such as entrepreneurial, computational, or design thinking, may also be relevant to their context. We resonate with their goal: to inspire further exploration of different ways of thinking in diverse contexts.
@@
(3) Values Thinking encompasses fundamental beliefs about right and wrong, and emphasizes ethical and normative considerations. It is about recognizing how decisions are made rather than determining their correctness, highlighting the importance of ethics, equity, and social justice within diverse cultural contexts. Values thinking may be viewed as self-reflection or introspection on one’s activities and assumptions, recognizing that perspectives may vary widely [11, p.110]. If students design an experiment to address nuclear waste, they can engage in a values thinking exercise by reflecting on the ethical implications of their proposed solutions. They may be encouraged to consider questions such as, “Who may be affected by our solution”. This process underscores the importance of identifying stakeholders, promoting a deeper understanding of the social and ethical context surrounding their scientific work.
@@ 4.2 English’s Framework
(1) Critical Thinking: Several undergraduate STEM reform documents emphasize critical thinking as essential for developing and refining ideas, whether through explanations, designs, or investigations. Engaging in argumentation and critique foster this thinking and deepen understanding [44, p.125],  [43, p.xv], [79, p.46]. Given this emphasis in reform documents, critical thinking rightly finds a place in English’s framework.
@@
While research indicates that argumentation can enhance conceptual learning [124, p.1304, p.1305][125], Osborne expresses concern about its notable absence in science education. He highlights the need for a deeper understanding of how argumentation facilitates learning and the characteristics of learning environments that foster effective student arguments [57]. In our exploration of SWoT frameworks within scientific argumentation and evidence-based reasoning, we found Slavit et al.’s application of Toulmin’s Claim-Evidence-Reasoning (CER) framework [126, 127, 125] in the context of integrated STEM to be particularly insightful [24, p.135][57, p.464]. An examination of Slavit et al.’s analytic coding framework provides valuable insights with potential applications for future qualitative analyses, particularly in applying CER within integrated STEM contexts.
@@
(2) Critical Mathematical Modelling: We were intrigued by English’s preference for the term ‘critical mathematical modelling’ over simply ‘mathematical modelling’. She defines modeling as the development of conceptual innovations that address real-world needs, emphasizing that effective modeling must include contextual understanding and critical analysis. Moreover, she argues that critiquing the outcomes of a model and reflecting on the lessons learned is crucial when addressing real-world problems [14, p.1222].
@@
Our laboratory work emphasizes group collaboration, and a significant portion of our research data is derived from group discussions. We appreciate English’s assertion that philosophical inquiry facilitates group problem-solving, peer-sharing of created models, and constructive feedback. In a recent computational physics laboratory activity, our students analyzed actual ‘big data’ from a High Energy Physics research laboratory [130]. Inspired by English’s framework, in our next iteration, we see scope to guide our students to pose critical questions, develop models, and engage in meaningful debates about their conclusions on the same problem.
@@
While our framework already incorporates metacognitive reflection as a way of thinking, we recognize that engaging in reflective thinking and ‘free-flowing’ [132, 21] conversations can further enhance philosophical inquiry. This integration could deepen students’ understanding and foster a richer educational experience.
@@ 5 Summary
In response to RG1, we introduced our WoT4EDP framework (see Figure 1) as a part of an ongoing and evolving discussion on STEM Ways of Thinking frameworks. In sections 3.1 and 3.2, we detailed the context of the Engineering Design-based physics problems within an introductory physics laboratory. Section 3.3 outlined our prior qualitative studies (see Table 1), placing us in a strong position to contribute to the literature on SWoT frameworks. Notably, we introduced computational thinking as a crucial element missing in other frameworks. Section 3.4 characterized the five core elements of our framework, supported by reform documents and relevant literature. Table 2 provided a basic operational scheme for applying the framework, emphasizing flexibility for practitioners. We believe that designing classroom instruction with a deliberate focus on the SWoT elements can help achieve diverse pedagogical goals, presented in a flexible and non-exhaustive list inspired by our experience as educators and supported by reform documents (see section 3.5). We acknowledged the limitations of our comparative study in section 3.6 and discussed the practical implications of our framework, highlighting simple yet practical benefits for students, teachers, and educators in section 3.7.
@@ 6 Future Work
A key challenge in qualitative research is the issue of overlapping codes, which we addressed extensively in our recently accepted manuscript. A common approach to qualitative analysis involves segmenting textual data, yet given the complexity of students’ reasoning and expression, multiple ways of thinking often co-occur within a single statement. While we previously illustrated such code co-occurrences with examples, qualitative analysis extends beyond mere segmentation. Over-segmentation, in particular, risks fragmenting students’ thought processes, raising critical questions about how best to delineate and interpret their ideas. We explored these issues in depth in our prior work and will expand on this discussion in the forthcoming study [21, p.7-10].
@@ 7 Acknowledgments
ChatGPT-4o and Perplexity were employed by the lead author to ‘wordsmith’ passages, and as a ‘learning partner’. We gratefully acknowledge the insightful and encouraging comments of anonymous referees on our earlier publications, which contributed to what we believe is a more coherent, detailed, and well-structured paper. We owe our gratitude to all the authors referenced, for they provided us with innumerable moments of much needed inspiration. A special thanks to Amir Bralin for designing the learning materials, with particular emphasis on scaffolding the Python-based activities and preparing the Jupyter Notebooks for the laboratory reports. We deeply appreciate the highly thoughtful insights Dr. Amogh Sirnoorkar offered towards the preparation of this manuscript. The lead author takes full responsibility for any error or omission.

### v4

@@ 3.7 Implications of our Framework
In spite of the aforementioned limitations, our WoT4EDP framework has significant implications for teaching and learning in STEM education, as well as educational research. By explicitly identifying and characterizing different forms of thinking—design, science, mathematical, metacognitive, and computational—this framework provides educators with a structured yet flexible approach to supporting and analyzing student thinking during engineering design-based tasks. For instance, instructors can use the framework to design activities that intentionally foster these thinking skills or assess student work with a focus on how well students integrate design and science thinking. The framework also offers guidance (see Enhancement approaches in Table 2) for scaffolding learning [80, 18], enabling educators to better identify gaps in student understanding and adapt instruction accordingly.
@@ 4 WoT Frameworks - A Comparative Study
FAWTEER emphasizes four specific Ways of Thinking: Futures, Values, Systems, and Strategic thinking. It conceptualizes Way of Thinking as a framework that can be practically applied to drive innovation and enhance researchers’ capacity to effect systemic change. Notably, Dalal et al. acknowledge that other ways of thinking, such as entrepreneurial, computational, or design thinking, may also be relevant to their context. We resonate with their goal: to inspire further exploration of different ways of thinking in diverse contexts.
@@
(3) Values Thinking encompasses fundamental beliefs about right and wrong, and emphasizes ethical and normative considerations. It is about recognizing how decisions are made rather than determining their correctness, highlighting the importance of ethics, equity, and social justice within diverse cultural contexts. Values thinking may be viewed as self-reflection or introspection on one’s activities and assumptions, recognizing that perspectives may vary widely [11, p.110]. If students design an experiment to address nuclear waste, they can engage in a values thinking exercise by reflecting on the ethical implications of their proposed solutions. They may be encouraged to consider questions such as, “Who may be affected by our solution?”. This process underscores the importance of identifying stakeholders, promoting a deeper understanding of the social and ethical context surrounding their scientific work.
@@ 4.2 English’s Framework
(1) Critical Thinking: Several undergraduate STEM reform documents emphasize critical thinking as essential for developing and refining ideas, whether through explanations, designs, or investigations. Engaging in argumentation and critique foster this thinking and deepen understanding [25, p.125],  [23, p.xv], [24, p.46]. Given this emphasis in reform documents, critical thinking rightly finds a place in English’s framework.
@@
While research indicates that argumentation can enhance conceptual learning [124, p.1304, p.1305][125], Osborne expresses concern about its notable absence in science education. He highlights the need for a deeper understanding of how argumentation facilitates learning and the characteristics of learning environments that foster effective student arguments [60]. In our exploration of SWoT frameworks within scientific argumentation and evidence-based reasoning, we found Slavit et al.’s application of Toulmin’s Claim-Evidence-Reasoning (CER) framework [126, 127, 125] in the context of integrated STEM to be particularly insightful [30, p.135][60, p.464]. An examination of Slavit et al.’s analytic coding framework provides valuable insights with potential applications for future qualitative analyses, particularly in applying CER within integrated STEM contexts.
@@
(2) Critical Mathematical Modelling: We were intrigued by English’s preference for the term critical mathematical modelling over simply mathematical modelling. She defines modeling as the development of conceptual innovations that address real-world needs, emphasizing that effective modeling must include contextual understanding and critical analysis. Moreover, she argues that critiquing the outcomes of a model and reflecting on the lessons learned is crucial when addressing real-world problems [14, p.1222].
@@
Our laboratory work emphasizes group collaboration, and a significant portion of our research data is derived from group discussions. We appreciate English’s assertion that philosophical inquiry facilitates group problem-solving, peer-sharing of created models, and constructive feedback. In a recent computational physics laboratory activity, our students analyzed actual big data from a High Energy Physics research laboratory [130]. Inspired by English’s framework, in our next iteration, we see scope to guide our students to pose critical questions, develop models, and engage in meaningful debates about their conclusions on the same problem.
@@
While our framework already incorporates metacognitive reflection as a way of thinking, we recognize that engaging in reflective thinking and “free-flowing” [132, 21] conversations can further enhance philosophical inquiry. This integration could deepen students’ understanding and foster a richer educational experience.
@@ 5 Summary
In response to RG1, we introduced our WoT4EDP framework (see Figure 1) as a part of an ongoing and evolving discussion on STEM Ways of Thinking SWoT frameworks. In sections 3.1 and 3.2, we detailed the context of the Engineering Design-based physics problems within an introductory physics laboratory. Section 3.3 outlined our prior qualitative studies (see Table 1), placing us in a strong position to contribute to the literature on SWoT frameworks. Notably, we introduced computational thinking as a crucial element missing in other frameworks. Section 3.4 characterized the five core elements of our framework, supported by reform documents and relevant literature. Table 2 provided a basic operational scheme for applying the framework, emphasizing flexibility for practitioners. We believe that designing classroom instruction with a deliberate focus on the SWoT elements can help achieve diverse pedagogical goals, presented in a flexible and non-exhaustive list inspired by our experience as educators and supported by reform documents (see section 3.5). We acknowledged the limitations of our comparative study in section 3.6 and discussed the practical implications of our framework, highlighting simple yet practical benefits for students, teachers, and educators in section 3.7.
@@ 6 Future Work
A key challenge in qualitative research is the issue of overlapping codes, which we addressed extensively in our recent paper [21]. A common approach to qualitative analysis involves segmenting textual data, yet given the complexity of students’ reasoning and expression, multiple ways of thinking often co-occur within a single statement. While we previously illustrated such code co-occurrences with examples, qualitative analysis extends beyond mere segmentation. Over-segmentation, in particular, risks fragmenting students’ thought processes, raising critical questions about how best to delineate and interpret their ideas. We explored these issues in depth in our prior work and will expand on this discussion in the forthcoming study [21, p.7-10].
@@ 7 Acknowledgments
ChatGPT-4o and Perplexity were employed by the lead author to wordsmith passages, and as a learning partner. We gratefully acknowledge the insightful and encouraging comments of anonymous referees on our earlier publications and the current manuscript, which contributed to what we believe is a more coherent, detailed, and well-structured paper. We owe our gratitude to all the authors referenced, for they provided us with innumerable moments of much needed inspiration. A special thanks to Amir Bralin for designing the learning materials, with particular emphasis on scaffolding the Python-based activities and preparing the Jupyter Notebooks for the laboratory reports. We deeply appreciate the highly thoughtful insights Dr. Amogh Sirnoorkar offered towards the preparation of this manuscript. The lead author takes full responsibility for any error or omission.